The temperature (T) of one mole of an ideal gas varies with its volume (V) as T = –V3 + V2, where  and
 are positive constants. The maximum pressure of gas during this process is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From
PV = RT
we get T = PV/R put in  T = –V3 + V2,
 PV/R = -aV^3 + bV^2
P = R(-aV^2 + bV)
dP/dV = 0 = R(-2aV + b) = 0
V = b/2a
Put in Pressure value we get,
P = R(-a*b^2/4a^2 + b*b/2a) = > R(b^2/2a – b^2/4a)